OpenStudy (anonymous):

Ted needs an average of at least 70 on his three history tests. He has already scored 85 and 60 on two tests. What is the minimum grade Ted needs on his third test?

OpenStudy (anonymous):

. Let t = Ted's score (85 + 60 + t)/3 ≥ 70 145 + t ≥ 210 t ≥ 65

OpenStudy (anonymous):

only 65 to get exactly 70

OpenStudy (anonymous):

85+65+60=210/3=70

OpenStudy (anonymous):

or [x+60+85]/3=70 if you solve for x=65

OpenStudy (anonymous):

(80+60+t)/3>70 t=210-140 t=70, can't get equal to sign